# GENESIS: A Generative Model of Episodic–Semantic Interaction

### v1

@@ 1 Introduction
A key distinction in psychology and neuroscience is between semantic and episodic memory, the two major forms of declarative memory. Semantic memory refers to structured knowledge about facts and concepts, primarily supported by cortical systems, whereas episodic memory involves personally experienced events embedded in specific spatial and temporal contexts, typically associated with the hippocampal formation (Tulving et al.,, 1972; Tulving,, 2002; Knowlton and Squire,, 1995).
@@
Despite this distinction, a long-held assumption is that semantic and episodic systems play complementary roles, as formalized in the influential complementary learning systems (CLS) theory (McClelland et al.,, 1995; Norman and O’Reilly,, 2003). The CLS theory posits that experiences are rapidly encoded in the hippocampus and later replayed to gradually train cortical semantic representations. Debate persists, however, over whether episodic traces remain permanently stored in the hippocampus after cortical consolidation. Recent proposals extend CLS theory while maintaining its central tenets. These include the ideas that hippocampal replay is goal-directed, supports statistical learning, and that cortical learning can sometimes be rapid rather than incremental (Kumaran et al.,, 2016; Kumaran and McClelland,, 2012; Schapiro et al.,, 2017). Another model that builds upon CLS introduces a Hopfield network for fast episodic encoding, coupled with a generative cortical model for semantic inference, allowing interactions such as pattern completion (Spens and Burgess,, 2024). Finally, a few models move beyond the standard assumptions of CLS. One focuses on the generative aspects of episodic memory and on semantic completion, whereby the hippocampus retrieves partial cues and cortical networks fill in missing semantic details (Fayyaz et al.,, 2022). Another theoretical proposal suggests that semantic memory engages in statistical learning, while episodic memory plays the complementary role of preserving surprising experiences in a relatively raw format (Nagy et al.,, 2025).
@@
Despite these advances, major challenges remain. A fundamental assumption in most existing models is that episodic memories are formed independently of semantic memories, typically within two separate networks (but see (Fayyaz et al.,, 2022; Nagy et al.,, 2025; Eleanor Spens,, 2025)). However, several converging lines of evidence suggest a more integrated picture. First, the hippocampus receives highly processed inputs from the entorhinal cortex, implying that episodic encoding inherently depends on cortical representations. Second, developmental findings show that semantic generalization can precede the formation of detailed episodic memories (Keresztes et al.,, 2018). Third, studies of narrative perception reveal that event representations are constructed within hierarchical cortical structures (Baldassano et al.,, 2017; Chang et al.,, 2022). Fourth, from a computational standpoint, processing complex stimuli such as images or words requires sophisticated cortical processing, making it implausible that episodic encoding could bypass these processes. A related limitation of most existing models is their assumption of a strictly unidirectional relationship between episodic and semantic memory, where the former merely trains the latter. This view struggles to account for phenomena such as semantic intrusions and gist-based distortions of memories, which are ubiquitous in episodic memory tasks (Schacter et al.,, 2011).
@@
Most critically, existing frameworks fail to jointly explain the core empirical phenomena that characterize both semantic and episodic memory. Semantic memory underlies the generalization of learned associations (e.g., 3–red, 5–blue, 3–blue) to novel combinations (e.g., 5–red), reflecting mechanisms of statistical learning (Ngo et al.,, 2021). It is also sensitive to the frequency and typicality of stimuli and categories (McCloskey,, 1980; Rosch et al.,, 1976). By contrast, episodic memory supports the “mental search” of past experiences. Recognition and free-recall studies—such as judging whether an image was previously seen (Brady et al.,, 2008) or recalling words from lists (Kahana,, 2020)—reveal robust behavioral regularities, including recency and serial-order effects, as well as semantic intrusions during retrieval. Moreover, episodic memory enables the constructive recombination of past experiences to generate novel scenarios, as shown in research on constructive episodic simulation and episodic future thinking (Schacter et al.,, 2017, 2015).
@@
None of the existing models of semantic–episodic interaction can account for the full range of phenomena across both domains. Moreover, they typically lack key mechanisms—such as the separation of content and context, and cue-dependent retrieval—that are central to leading models of memory search (Kahana,, 2020; Norman et al.,, 2008; Atkinson and Shiffrin,, 1968; Howard and Kahana,, 2002; Farrell and Lewandowsky,, 2002; Sederberg et al.,, 2008; Raaijmakers and Shiffrin,, 1981; Gillund and Shiffrin,, 1984; Shiffrin and Steyvers,, 1997). As a result, we still lack an integrative computational framework capable of jointly explaining the main empirical findings on semantic and episodic memory and their dynamic interactions.
@@ 2.1 The Generative Episodic–Semantic Integration System (GENESIS) model
The Generative Episodic–Semantic Integration System (GENESIS) model is schematically illustrated in Figure 1. The model comprises two interconnected generative models (limited-capacity variational autoencoders or VAEs (Kingma and Welling,, 2013; Higgins et al.,, 2017)) – a Cortical-VAE and a Hippocampal-VAE (Figure 1) – along with an Episodic Memory component, which we implement for simplicity using a Retrieval-Augmented Generation (RAG) system (Lewis et al.,, 2020).
Figure 1: Illustration of the Generative Episodic–Semantic Integration System (GENESIS) model. A. The Cortical-VAE (a conditional $\beta$-VAE (Sohn et al.,, 2015)) achieves a representative input embedding by merging the Gaussian latent state $z$ with color and digit conditional embeddings, $e_{C},e_{S}$, respectively, and provides the input to either the cortical decoder (blue arrow) or the Hippocampal-VAE (red arrow). B. The Hippocampal-VAE (a $\beta$-VAE (Kingma and Welling,, 2013; Higgins et al.,, 2017)) generates a compressed representation of the input embedding coming from the Cortical-VAE encoding, and incorporates temporal embeddings, to form an episode-specific key. C. Episodic memory is implemented as a Retrieval-Augmented Generation (RAG) system that memorizes episodes as key-value pairs (or sequences of key-value pairs). Values are item embeddings (possibly compressed by the Cortical-VAE). Keys are a mixture of item embeddings (further compressed by the Hippocampal-VAE) and temporal embeddings. D. Interaction between cortical and hippocampal systems given a query. The query triggers a query-key-value search process that retrieves one or more episodes from Episodic Memory and routes them to the decoder of the Cortical-VAE. See the main text for further explanation.
@@
This latent embedding is then processed along two parallel pathways (blue and red arrows in Figure 1). First, it can be decoded by the Cortical-VAE’s decoder, corresponding to cortical (here, visual) reconstruction. Second, it can be routed to hippocampal circuits to form an episodic memory – which boils down to creating a key–value pair. To do so, the item embedding is further encoded and decoded through the (limited-capacity) Hippocampal-VAE and combined with a temporal embedding that captures when the item was experienced. The resulting compressed representation forms the key, while the associated value corresponds to the Cortical-VAE item embedding. These key–value pairs are stored in a retrieval-augmented memory (RAG) architecture (Lewis et al.,, 2020), where each pair constitutes an episode. More generally, an episode can also be viewed as a short sequence of such key–value pairs.
@@
As assumed by models of episodic search (Kahana,, 2020), queries initiating recall can arise from different sources in different psychological paradigms. In recognition (old/new) tasks, a probe image (e.g., a 3-yellow) is encoded by the Cortical-VAE; its embedding serves as the query. High query–key similarity indicates that the item was already known and low query–key similarity indicates that the item was new. In free or serial recall paradigms, recall is triggered by a temporal embedding corresponding to a specific time point, such as the moment when the participant is asked to recall. Existing models of episodic search sometimes extend the temporal embedding with a list identifier, but we do not use this mechanism in our simulations for simplicity. Retrieval can proceed iteratively, with the key of each recalled item serving as the query for the next, thereby generating a temporal sequence of remembered episodes.
@@ 2.2 Statistical learning, capacity limitations, and generalization in semantic memory
A central tenet of most accounts of semantic memory, including ours, is that it implements a form of statistical learning by gradually accumulating structured knowledge from experience—either directly perceived or replayed from episodic memory (McClelland et al.,, 1995; Schapiro et al.,, 2017). Here, we test the capacity of the Cortical-VAE to learn the statistical regularities of input images and to generalize by recombining learned elements in novel ways. Moreover, in line with evidence that semantic processing is limited by attention and memory (Marois and Ivanoff,, 2005; Zenon et al.,, 2019), we examine how encoding capacity in the Cortical-VAE shapes reconstruction quality and representational geometry.
@@
To implement the Cortical-VAE, we employ a convolutional conditional $\beta$-VAE with limited capacity (Higgins et al.,, 2017; Burgess et al.,, 2018; Sohn et al.,, 2015). This deep generative model includes an encoder, which compresses an input (e.g., an image) into a latent representation, and a decoder, which reconstructs the input from that latent representation (Figure 1A). The Cortical-VAE was trained in an unsupervised manner on colored MNIST digits to encode and reconstruct images. During encoding, each input image is mapped to a composite item embedding $z,e_{C},e_{S}]$, where $z$ is a Gaussian latent vector capturing item-specific variability, and $e_{C}$ and $e_{S}$ are discrete embeddings representing color and digit identity, respectively.
@@
Figure 2: Simulation of statistical learning, capacity limitations, and generalization in semantic memory. A. Theoretical rate-distortion curve (Shannon et al.,, 1959). The conditional $\beta$-VAE enables the implementation of the rate-distortion trade-off in the loss terms, as shown in the formula at the top ($e=[e_{S},e_{C}]$ is the set of conditioning embeddings) B. Reconstruction performance of images as a function of Cortical-VAE capacity levels C. Multidimensional Scaling (MDS) of the latent feature maps of images (only red digits for simplicity) processed by the Cortical-VAE under low-capacity (0 nats, left) and high-capacity (8 nats, right). Feature maps are extracted from the first decoder layer. D. Schematic of the Leave-One-(Pair)-Out (LOO) protocol. The Cortical-VAE is trained on an ablated dataset with one digit–color pair (e.g., 5-red) withheld; the missing pair is later generated by combining the corresponding conditional embeddings. E. Simulated exemplars for each left-out digit–color pair. Each grid cell shows examples generated from a model trained with that specific pair ablated. F. Heatmaps indicating the classification accuracy of the generated digit–color pairs. See the main text for further explanation.
@@
To account for limited capacity in the Cortical-VAE, we leverage the fact that the $\beta$ parameter of the $\beta$-VAE has a well-defined theoretical interpretation as a capacity bottleneck within rate–distortion theory (Shannon et al.,, 1959), where decreases in reconstruction accuracy with limited resources follow a characteristic rate–distortion curve (Figure 2A). This framework treats encoding capacity as a proxy for available cognitive resources (e.g., attention, memory) during learning, which may vary with cognitive demands (e.g., performing a dual task may reduce capacity) or contextual factors (e.g., salient or surprising events may recruit greater capacity) (Nagy et al.,, 2025; D’Amato et al.,, 2025; Bates and Jacobs,, 2020; Bhui et al.,, 2021; Sims,, 2016, 2003, 2018; Gershman,, 2021; Jakob and Gershman,, 2023; Anderson and Milson,, 1989; Pezzulo et al.,, 2024).
@@
To systematically examine how limited encoding capacity in the Cortical-VAE affects reconstruction quality, we trained the model under varying capacity constraints (0,1,2,4,8{0,1,2,4,8} nats). As expected, the model trained at high capacity accurately reconstructed MNIST images. This finding aligns with extensive prior work showing that $\beta$-VAEs support robust statistical learning of visual stimuli, whether trained on real inputs or through generative replay (Higgins et al.,, 2017; Burgess et al.,, 2018; Shin et al.,, 2017; Van de Ven et al.,, 2020; Stoianov et al.,, 2022). Reducing capacity led to a progressive decline in reconstruction accuracy, following the characteristic rate–distortion curve of rate–distortion theory (Figure 2B). Interestingly, high- and low-capacity training also produced distinct latent geometries. Figure 2B–C compares the first-layer decoder representations at high capacity (8 nats, right) and low capacity (0 nats, left). At high capacity, representations of different instances of the same digit (colored dots) spanned a wide region of latent space, whereas at low capacity they became less differentiated and collapsed into a compact region, with all instances converging toward a single prototype representation.

### v2

@@ 1 Introduction
A key distinction in psychology and neuroscience is between semantic and episodic memory, the two major forms of declarative memory. Semantic memory refers to structured knowledge about facts and concepts, primarily supported by cortical systems, whereas episodic memory involves personally experienced events embedded in specific spatial and temporal contexts, mainly associated with the hippocampal formation (Tulving and others, 1972; Tulving, 2002; Knowlton and Squire, 1995; Eichenbaum, 1999).
@@
Despite this distinction, a long-held assumption is that semantic and episodic systems play complementary roles, as formalized in the influential complementary learning systems (CLS) theory (McClelland et al., 1995; Norman and O’Reilly, 2003). The CLS theory posits that experiences are rapidly encoded in the hippocampus and later replayed to gradually train cortical semantic representations. Debate persists, however, over whether episodic traces remain permanently stored in the hippocampus after cortical consolidation. Recent proposals extend CLS theory while maintaining its central tenets. These include the ideas that hippocampal replay is goal-directed, supports statistical learning, and that cortical learning can sometimes be rapid rather than incremental (Kumaran et al., 2016; Kumaran and McClelland, 2012; Schapiro et al., 2017). Another model that builds upon CLS introduces a Hopfield network for fast episodic encoding, coupled with a generative cortical model for semantic inference, allowing interactions such as pattern completion (Spens and Burgess, 2024). Finally, a few models move beyond the standard assumptions of CLS. One focuses on the generative aspects of episodic memory and on semantic completion, whereby the hippocampus retrieves partial cues and cortical networks fill in missing semantic details (Fayyaz et al., 2022). Another theoretical proposal suggests that semantic memory engages in statistical learning, while episodic memory plays the complementary role of preserving surprising experiences in a relatively raw format (Nagy et al., 2025).
@@
Despite these advances, major challenges remain. Most existing models assume that episodic memories are formed independently of semantic memories, by implementing them within two separate networks (but see (Fayyaz et al., 2022; Nagy et al., 2025; Spens and Burgess, 2025)). However, several converging lines of evidence suggest a more integrated picture. First, the hippocampus receives highly processed inputs from the entorhinal cortex, implying that episodic encoding inherently depends on cortical representations. Second, developmental findings show that semantic generalization can precede the formation of detailed episodic memories (Keresztes et al., 2018). Third, studies of narrative perception reveal that event representations are constructed within hierarchical cortical structures (Baldassano et al., 2017; Chang et al., 2022). Fourth, from a computational standpoint, processing complex stimuli such as images or words requires sophisticated cortical processing, making it implausible that episodic encoding could bypass these processes. A related limitation of most existing models is their assumption of a strictly unidirectional relationship between episodic and semantic memory, where the former merely trains the latter. This view struggles to account for phenomena such as semantic intrusions and gist-based distortions of memories, which are ubiquitous in episodic memory tasks (Schacter et al., 2011).
@@
Most critically, existing frameworks fail to jointly explain the core empirical phenomena that characterize both semantic and episodic memory. Semantic memory underlies the generalization of learned associations (e.g., 3–red, 5–blue, 3–blue) to novel combinations (e.g., 5–red), reflecting mechanisms of statistical learning (Ngo et al., 2021). It is also sensitive to the frequency and typicality of stimuli and categories (McCloskey, 1980; Rosch et al., 1976). By contrast, episodic memory supports the “mental search” of past experiences. Recognition and free-recall studies—such as judging whether an image was previously seen (Brady et al., 2008) or recalling words from lists (Kahana, 2020)—reveal robust behavioral regularities, including recency and serial-order effects, as well as semantic intrusions during retrieval. Moreover, episodic memory enables the constructive recombination of past experiences to generate novel scenarios, as shown in research on constructive episodic simulation and episodic future thinking (Schacter et al., 2017, 2015).
@@
None of the existing models of semantic–episodic interaction can account for the full range of phenomena across both domains. Moreover, they typically lack key mechanisms—such as the separation of content and context, and cue-dependent retrieval—that are central to leading models of memory search (Kahana, 2020; Norman et al., 2008; Atkinson and Shiffrin, 1968; Howard and Kahana, 2002; Farrell and Lewandowsky, 2002; Sederberg et al., 2008; Raaijmakers and Shiffrin, 1981; Gillund and Shiffrin, 1984; Shiffrin and Steyvers, 1997; Nelson and Shiffrin, 2013; Sederberg et al., 2011). As a result, we still lack an integrative computational framework capable of jointly explaining the main empirical findings on semantic and episodic memory and their dynamic interactions.
@@ 2.1 The Generative Episodic–Semantic Integration System (GENESIS) model
The Generative Episodic–Semantic Integration System (GENESIS) model is schematically illustrated in Figure 1. The model comprises two interconnected generative models (limited-capacity variational autoencoders or VAEs (Kingma and Welling, 2013; Higgins et al., 2017)) – a Cortical-VAE and a Hippocampal-VAE (Figure 1) – along with an Episodic Memory component, which we implement for simplicity using a Retrieval-Augmented Generation (RAG) system (Lewis et al., 2020).
Figure 1: Illustration of the Generative Episodic–Semantic Integration System (GENESIS) model. The violet and yellow boxes denote core components of the semantic and episodic systems, respectively. Although shown separately, these systems closely interact, as illustrated in the white panel. A. The Cortical-VAE (a conditional $\beta$-VAE (Sohn et al., 2015)) achieves a representative input embedding by merging the Gaussian latent state $z$ with color and digit conditional embeddings, $e_{C},e_{S}$, respectively, and provides the input to either the cortical decoder (blue arrow) or the Hippocampal-VAE (red arrow). B. The Hippocampal-VAE (a $\beta$-VAE (Kingma and Welling, 2013; Higgins et al., 2017)) generates a compressed representation of the input embedding coming from the Cortical-VAE encoding, and incorporates temporal embeddings, to form an episode-specific key. C. Episodic memory is implemented as a Retrieval-Augmented Generation (RAG) system that memorizes episodes as key-value pairs (or sequences of key-value pairs). Values are item embeddings (possibly compressed by the Cortical-VAE). Keys are a mixture of item embeddings (further compressed by the Hippocampal-VAE) and temporal embeddings. D. Interaction between cortical and hippocampal systems given a query. The query triggers a query-key-value search process that retrieves one or more episodes from Episodic Memory and routes them to the decoder of the Cortical-VAE. See the main text for further explanation.
@@
This latent embedding is then processed along two parallel pathways (blue and red arrows in Figure 1). First, it can be decoded by the Cortical-VAE’s decoder, corresponding to cortical (here, visual) reconstruction. Second, it can be routed to hippocampal circuits to form an episodic memory – which boils down to creating a key–value pair. To do so, the item embedding is further encoded and decoded through the (limited-capacity) Hippocampal-VAE and combined with a temporal embedding that captures when the item was experienced. The resulting compressed representation forms the key, while the associated value corresponds to the Cortical-VAE item embedding. These key–value pairs are stored in a retrieval-augmented memory (RAG) architecture (Lewis et al., 2020), where each pair constitutes an episode. More generally, an episode can also be viewed as a short sequence of such key–value pairs.
@@
As assumed by models of episodic search (Kahana, 2020), queries initiating recall can arise from different sources in different psychological paradigms. In recognition (old/new) tasks, a probe image (e.g., a 3-yellow) is encoded by the Cortical-VAE; its embedding serves as the query. High query–key similarity indicates that the item was already known and low query–key similarity indicates that the item was new. In free or serial recall paradigms, recall is triggered by a temporal embedding corresponding to a specific time point, such as the moment when the participant is asked to recall. Existing models of episodic search sometimes extend the temporal embedding with a list identifier, but we do not use this mechanism in our simulations for simplicity. Retrieval can proceed iteratively, with the key of each recalled item serving as the query for the next, thereby generating a temporal sequence of remembered episodes.
In summary, although Figure 1 depicts the core components of the episodic system (yellow boxes) as distinct from those of the semantic system (violet box), both episodic memory formation and recall arise from their interaction. Specifically, they depend on the Cortical-VAE (providing encoding and decoding functions), the Hippocampal-VAE (generating keys), and the RAG module (supporting query–key–value matching). Accordingly, semantic and episodic processes are not strictly separated or localized within isolated modules, but emerge from coordinated interactions among these components.
@@ 2.2 Statistical learning, capacity limitations, and generalization in semantic memory
A central tenet of most accounts of semantic memory, including ours, is that it implements a form of statistical learning by gradually accumulating structured knowledge from experience—either directly perceived or replayed from episodic memory (McClelland et al., 1995; Schapiro et al., 2017). Here, we test the capacity of the Cortical-VAE to learn the statistical regularities of input images and to generalize by recombining learned elements in novel ways. Moreover, in line with evidence that semantic processing is limited by attention and memory (Marois and Ivanoff, 2005; Zenon et al., 2019), we examine how encoding capacity in the Cortical-VAE shapes reconstruction quality and representational geometry.
@@
To implement the Cortical-VAE, we employ a convolutional conditional $\beta$-VAE with limited capacity (Higgins et al., 2017; Burgess et al., 2018; Sohn et al., 2015). This deep generative model includes an encoder, which compresses an input (e.g., an image) into a latent representation, and a decoder, which reconstructs the input from that latent representation (Figure 1A). The Cortical-VAE was trained in an unsupervised manner on colored MNIST digits to encode and reconstruct images. During encoding, each input image is mapped to a composite item embedding $z,e_{C},e_{S}]$, where $z$ is a Gaussian latent vector capturing item-specific variability, and $e_{C}$ and $e_{S}$ are discrete embeddings representing color and digit identity, respectively.
@@
Figure 2: Simulation of statistical learning, capacity limitations, and generalization in semantic memory. A. Theoretical rate-distortion curve (Shannon and others, 1959). The conditional $\beta$-VAE enables the implementation of the rate-distortion trade-off in the loss terms, as shown in the formula at the top ($e=[e_{S},e_{C}]$ is the set of conditioning embeddings) B. Reconstruction performance of images as a function of Cortical-VAE capacity levels C. Multidimensional Scaling (MDS) of the latent feature maps of images (only red digits for simplicity) processed by the Cortical-VAE under low-capacity (0 nats, left) and high-capacity (8 nats, right). Feature maps are extracted from the first decoder layer. D. Schematic of the Leave-One-(Pair)-Out (LOO) protocol. The Cortical-VAE is trained on an ablated dataset with one digit–color pair (e.g., 5-red) withheld; the missing pair is later generated by combining the corresponding conditional embeddings. E. Simulated exemplars for each left-out digit–color pair. Each grid cell shows examples generated from a model trained with that specific pair ablated. F. Heatmaps indicating the classification accuracy of the generated digit–color pairs. See the main text for further explanation.
@@
To account for limited capacity in the Cortical-VAE, we leverage the fact that the $\beta$ parameter of the $\beta$-VAE has a well-defined theoretical interpretation as a capacity bottleneck within rate–distortion theory (Shannon and others, 1959), where decreases in reconstruction accuracy with limited resources follow a characteristic rate–distortion curve (Figure 2A). This framework treats encoding capacity as a proxy for available cognitive resources (e.g., attention, memory) during learning, which may vary with cognitive demands (e.g., performing a dual task may reduce capacity) or contextual factors (e.g., salient or surprising events may recruit greater capacity) (Nagy et al., 2025; D’Amato et al., 2025; Bates and Jacobs, 2020; Bhui et al., 2021; Sims, 2016, 2003, 2018; Gershman, 2021; Jakob and Gershman, 2023; Anderson and Milson, 1989; Pezzulo et al., 2024).
@@
To systematically examine how limited encoding capacity in the Cortical-VAE affects reconstruction quality, we trained the model under varying capacity constraints (0,1,2,4,8{0,1,2,4,8} nats). As expected, the model trained at high capacity accurately reconstructed MNIST images. This finding aligns with extensive prior work showing that $\beta$-VAEs support robust statistical learning of visual stimuli, whether trained on real inputs or through generative replay (Higgins et al., 2017; Burgess et al., 2018; Shin et al., 2017; Van de Ven et al., 2020; Stoianov et al., 2022). Reducing capacity led to a progressive decline in reconstruction accuracy, following the characteristic rate–distortion curve of rate–distortion theory (Figure 2B). Interestingly, high- and low-capacity training also produced distinct latent geometries. Figure 2B–C compares the first-layer decoder representations at high capacity (8 nats, right) and low capacity (0 nats, left). At high capacity, representations of different instances of the same digit (colored dots) spanned a wide region of latent space, whereas at low capacity they became less differentiated and collapsed into a compact region, with all instances converging toward a single prototype representation.